Operations Report – April 16th

Crew 297 Operations Report 16-04-2024

SOL: 2

Name of person filing report: Matt Lynch (Engineer)

Non-nominal systems: Suit 2 (stuck valve) and Suit 9 (broken visor).

Notes on non-nominal systems: The issues with Suit 2 and 9 were already noted before our mission. We have fixed the issues with Suit 3 today, thanks to the advice of Scott Davis. Suit’s 4 issues were investigated and likely fixed, however further observation will be done in the coming days to confirm this.


Opportunity rover used:

Hours: 178.8

Beginning charge: 100%

Ending charge: 63%

Currently charging: No

Curiosity rover used:

Hours: 277.5

Beginning charge: 100%

Ending charge: 53%

Currently charging: No

General notes on rovers: Some tape on Curiosity’s chassis has peeled back and is sticking to the driver’s suit, this should be replaced with new tape on the next EVA.

Summary of Hab operations:

WATER USE: 20.1 gallons

Water (static tank): 400.4 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: yes

Summary of internet: Looking up recipes, downloading entertainment, downloading navigation applications, report formatting methods, catching up on Earthly news

Summary of suits and radios:

Radios 2,3,4, and 6 were used.

EVA Suits 1,3, and 6 were used.

Summary of GreenHab operations: Watered and first time harvesting some tomatoes today

WATER USE: 2.0 gallons

Heater: On

Supplemental light: N/A

Harvest: 49g of cherry tomatoes

Summary of ScienceDome operations: NA

Summary of RAM operations: (Tools used, work done): A drill and snips were used last night to prepare the NPS for the EVA today. Several screwdrivers and wrenches were used in the repairing of Suits 3 & 4 (soldering was done without a crewmember’s personal tools).

Summary of any observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site